OpenStudy (kweenlexis):

What is the slope of the line? ​ 20x−5y=12 Enter your answer in the box.

OpenStudy (kweenlexis):

4?

OpenStudy (will.h):

20x -5y = 12 -5y = -20x + 12 Y = 4x -12/5 So yeah 4

OpenStudy (kweenlexis):

Thank u

OpenStudy (will.h):

U welcome